The mass spectrum of an organic compound shows the relative abundances of M to be 53.76% and M 1 to be 3.489%. Assuming the peaks are caused by 12C and 13C isotopes, determine the number of carbon atoms in the compound. Natural abundances: 12C is 98.93% and 13C is 1.07%
Answer #1

Solution:

According to relative abundance and natural abundance equation,

AM+1 / AM = nc x ( abundance 13C / abundance 12C)

Where, AM+1= Relative percent abundance of M+1 peak

AM = Relative percentage abundance of M peak

Therefore,

AM1 / AM = nc x (1.07/98.93)

3.489 / 53.76 = nc x 1.07 / 98.93

6.4205 = nc x 1.07

nc = 6.4205 / 1.07 = 6.00

Hence, number of carbons = 6